OpenStudy (anonymous):

help simplify- with work shown please :) (5i)(3-2i)

OpenStudy (anonymous):

*remember \[i^2= -1\] and when you distribute, you get \[15i - 10i^2\] so...?

OpenStudy (anonymous):

im still lost :(

OpenStudy (anonymous):

Do you know that \[i = \sqrt{-1}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is the first step to dist. 5i to 3-2i ?

OpenStudy (anonymous):

? so \[i^2 = -1\] If you distribute the 5i to the numbers inside the ( )

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

from 15i-10i^2 i dont know what to do! :(

OpenStudy (anonymous):

(5i)(3-2i) =(5i)*(3-2i) =5i*3+5i*(-2i) =15i-10i^2 =15i-10*(-1) <-- i^2=-1 =15i+10